Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 596
Entry time: 25 Jan 2019, 10:07
|
Author: |
Adriano Violante |
Location: |
General |
Severity: |
Information |
Subject: |
Report after dispatch meeting 09.00 |
|
|
SPB/SFX
- Trigger for Jungfrau detector is needed.
- Commissioning of the instrument should start next week
- Loop priorities for the donwstream components have been set with AE.
FXE
SQS
- VAC interlock test will be done during next days after update of the loop.
- Loop 4 has to be updated following some HW changes. Should run by next week since is needed for the preparation of the users run.
- Digitizers crate is being tested by AE and CAS and will be installed/usable by the middle of next week.
- Burst trigger - status and requirements will be discussed with BFernandes after this meeting.
- MPOT test CY/TB Today at 13:30
SCS
MID
HED
- Preparation of Hi res mono has been completed last week
- OPT hutch CRL commissioning ongoing. ATT will follow (needed missing crate will be installed next week)
- Chamber motion for the monochromator has to be checked
OPT Lasers
XRO
- Loops update in thetunnels have been completed. the MOV1 in SASE3 included also some PBLMs ==> XRD should check whether everything is fine afte the update.
XPD
VAC
- Tunnels are ready for beam.
AE
- EXAPOD in SPB can be tested.
- HED need PPU in XTD6. To be re-checked.
- The timing software needs to be updated. BFernandes circulated a mail with proposed date: Next monday at 14:00. In case of different suggestion, please contact Bruno. CAS suggests to keep the proposed date.
ITDM
ET
CAS
- Data logger schould be working for every component now. ==> Meeting later with VAC and Laser group to define the use case and start a new optimization project.
- VAC of HED Opt tests are ongoing (CY) before pumping down.
- For the debug of gauges control is possible that a ZZ access will be requested. (CY)
AoB
- Tunnel access schedule will be updated by HSinn after the Operation meeting with the latest information.
|